Given f(x) = –5x + 3, find f(a + 1).
Found 2 solutions by jim_thompson5910, Alan3354:
Answer by jim_thompson5910(35256)

Wherever you see 'x' replace it with 'a+1'


f(x) = -5x + 3


f(a+1) = -5(a+1) + 3


f(a+1) = -5a - 5 + 3


f(a+1) = -5a - 2

Answer by Alan3354(69443)

sub a+1 for x